# Understanding Ranking LLMs: A Mechanistic Analysis for Information Retrieval

### v2

@@ Abstract.
Transformer networks, particularly those achieving performance comparable to GPT models, are well known for their robust feature extraction abilities. However, the nature of these extracted features and their alignment with human-engineered ones remain unexplored. In this work, we investigate the internal mechanisms of state-of-the-art, fine-tuned LLMs for passage reranking. We employ a probing-based analysis to examine neuron activations in ranking LLMs, identifying the presence of known human-engineered and semantic features. Our study spans a broad range of feature categories, including lexical signals, document structure, query-document interactions, and complex semantic representations, to uncover underlying patterns influencing ranking decisions.
Through experiments on four different ranking LLMs, we identify statistical IR features that are prominently encoded in LLM activations, as well as others that are notably missing. Furthermore, we analyze how these models respond to out-of-distribution queries and documents, revealing distinct generalization behaviors. By dissecting the latent representations within LLM activations, we aim to improve both the interpretability and effectiveness of ranking models. Our findings offer crucial insights for developing more transparent and reliable retrieval systems, and we release all necessary scripts and code to support further exploration.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1.1. Contributions and Findings
    We discovered that several human-engineered metrics, such as covered query term number, covered query term ratio, min of term frequency, min and mean of stream length normalized term frequency, and variance of tf*idf, are prominently encoded within LLM activations. In contrast, certain features, including sum of stream length normalized term frequency, max of stream length normalized term frequency, and BM25, show no discernible representation in LLM activations.
@@
    We fine-tuned and probed two additional reranking models, Llama3.1-8b and Pythia-6.9b, resulting in RankLlama3-8b and RankPythia, alongside probing RankLlama2-7b and RankLlama2-13b. The results revealed broadly consistent outcomes across all models, suggesting that different LLM architectures, when fine-tuned similarly, encode latent features in a comparable way within their activations.
@@
    Finally, we show further evidence on the presence of these encoded features using Shapley values. All scripts, datasets, and fine-tuned ranking LLMs developed and/or used in this study are made publicly available. [^1] https://suppressed.for.review
@@ 2.1. Mechanistic Interpetability & Sparse Probing
Mechanistic Interpretability is the ambitious goal of gaining an algorithmic-level understanding of any deep neural network’s computations (Räuker et al., 2023). This goal can be achieved in numerous ways, some of them being by studying weights (e.g., weight masking (Csordás et al., 2020; Wortsman et al., 2020), by continual learning (De-Arteaga et al., 2019)), by studying individual neurons (e.g., excitement based (Zhou et al., 2014), activation patching (Makelov et al., 2023), gradient-based (Ancona et al., 2019), perturbation and ablation based  (Zhou et al., 2018)), by studying subnetworks (e.g., sparsity based  (Meister et al., 2021), circuit analysis based  (Fiacco et al., 2019)), by studying representations (e.g., tokens  (Li et al., 2021), attention  (Clark et al., 2019), probing  (Belinkov, 2022; Gurnee and Tegmark, 2023; Chen et al., 2023)), by training sparse auto encoders  (Makelov et al., 2024; Kissane et al., 2024; Cunningham et al., 2023) etc. In this study, we focus on studying the activations of individual neurons and also some groups of neurons with the help of probing regressors. To that end, we use concepts from a host of techniques mentioned above.
@@
Probing aims to determine if a given representation effectively captures a specific type of information, as discussed by Belinkov (Belinkov, 2022). This technique employs transfer learning to test whether embeddings contain information pertinent to a target task. The three essential steps in probing include: (1) obtaining a dataset with examples that exhibit variation in a particular quality of interest, (2) embedding these examples, and (3) training a model on these embeddings to assess if it can learn the quality of interest. This method is versatile as it can utilize any inner representation from any model. However, a limitation of probing methods is that a successful probe does not necessarily mean that the probed model actually utilizes that information about the data  (Ravichander et al., 2020). Belinkov (Belinkov, 2022) provides a comprehensive survey on probing methods for large language models, discussing their advantages, disadvantages, and complexities. Gurnee et al.  (Gurnee et al., 2023) further introduce sparse probing, where they mine features of interest over groups of representations up to $k$ in size. We build upon sparse probing in this work.
@@
Table 1. The Ranking effectiveness of various fine-tuned ranking LLMs, in comparision to the performance of a BM25 reranker. We observe that while LLMs are effecient for reranking, they are opaque and complex to understand.
@@ 2.2. Inner Interpretability in IR
Most works of interpretability in information retrieval  (Anand et al., 2022) have been model extrinsic (Rahimi et al., 2021; Singh and Anand, 2019; Chowdhury et al., 2023). Among model intrinsic interpretability methods, some works focus on using gradient-based methods to identify important neurons  (Chen et al., 2024; Fernando et al., 2019). Although gradient based methods give an accurate perspective on the flow of information, they are too fine-grained to give a human-level understanding of the LLM’s inner circuit (Anand et al., 2022). A number of works have attempted to examine the inner workings of neural retrievers to understand if they satisfy IR axioms and/or to spot known features. Fan et al. (Fan et al., 2021) probe fine-tuned BERT models on three different tasks, namely document retrieval, answer retrieval, and passage retrieval, to find out if these different forms of relevance lead to different models. Zhan et al.  (Zhan et al., 2020) study the attention patterns of BERT after fine-tuning on the document ranking task and make note of how BERT dumps duplicate attention weights on high frequency tokens (like periods). Similar to them, Choi et al. (Choi et al., 2022) study attention maps in BERT and report the discovery of the IDF feature within BERT attention patterns. ColBERT investigations (Formal et al., 2021, 2022) study its term-matching mechanism and conclude that it captures a notion of term importance, which is enhanced by fine-tuning. MacAvaney et al.  (MacAvaney et al., 2022) propose ABNIRML a set of diagnostic probes for neural retrieval models that allow searching for features like writing styles, factuality, sensitivity and word order in models like ColBERT and T5. Recently Parry et al.  (Parry et al., 2024) provide a framework for diagnostic analysis and intervention within Neural Ranking models by generating perturbed inputs followed by activation patching.
These studies, however, have been conducted in the pre-LLM era and hence are much smaller in size. It is unknown if the findings from BERT will carry over to LLMs like Llama, given that BERT is an encoder-only model whereas most modern LLMs are decoder-only models. While most of the above works focus on attention heads and pattern, our work focuses on probing the MLP activation layers, which are now believed to be the primary feature extraction location within the LLM (Gurnee et al., 2023). To the best of our knowledge, this is one of the first works towards study of neuron activations in LLMs for a large scale of statistical and semantic features.
@@ 3. Identifying Context Neurons
Figure 2. Probing for statistical features from the MSLR dataset in RankLlama2-7b model. Here $QT$ stands for Query Term, $TF$ stands for Term Frequency and $\cdot/L$ stands for length normalized. The graph lines indicate the presence of a particular feature along the layers of the LLM. Certain features like $Min\ TF*IDF$ show consistent presence across the layers. Other features like $Covered\ QT\ Number$, $Covered\ QT\ Ratio$, $Mean(TF/L)$ and $Variance\ TF*IDF$ show increasing prominence from the 1st layer to the last, ultimately playing an important role in ranking decision making. Other MSLR features like $Sum(TF/L)$, $Max(TF/L)$, and $Sum\ TF*IDF$ show negative correlation with RankLlama decision making.
@@
Target Features: The MSLR dataset (Qin and Liu, 2013) provides a comprehensive list of features that have been historically recognized as highly effective for ranking models (Han and Lei, 2018; Hochma et al., 2024). Consequently, we aim to search for a subset of these MSLR features within the LLM activations. Specifically, we focus on mining the following 19 features from the MSLR dataset within the LLM (“stream” means the document or passage text): Min TF, Min(TF/L), Min TF*IDF, Covered QT Number, Covered QT Ratio, Mean(TF/L), Variance TF*IDF, BM25, Mean TF*IDF, Variance(TF/L), Mean TF, Variance TF, Sum TF, Max TF, Max(TF/L), Stream length, Sum TF, Max TF*IDF and Sum TF*IDF.
@@
In addition to MSLR features, we mine for known query-document similarity metrics within the LLM architecture. For this we consider traditional query-document similarity scores like tf*idf cosine scores, Euclidian score, Manhattan score, KL-divergence score, Jensen-Shannon divergence score as well as popular semantic relevance metrics like BERT and T5 scores.
@@
Each feature is probed individually; for instance, the probe for stream length is dedicated solely to that feature without searching for any others. Although many of these features are known to be correlated and may share neurons, this study does not include correlation analysis amongst features.
@@
Probing Datasets: To facilitate probing (Sajjad et al., 2022), we need to curate a dataset of input sequences to study the model. We select query-document pairs from the MS MARCO test set (Nguyen et al., 2016) for this purpose. For each query, we include documents that are highly relevant, highly irrelevant, and of intermediate relevance. Our input sequences for the ranking LLM consist of these query-document pairs. Note that it is important for each probing dataset to be balanced – i.e., contain a uniform number of samples across the range of values of the feature being probed (Gurnee et al., 2023; Belinkov, 2022). For instance, when analyzing the BM25 feature, our dataset included a balanced ratio of query-documents with both low and high BM25 values. Having an unbalanced probing dataset can lead to a biased analysis as the model might have a tendency to overfit the dominant class. For each input sequence, we calculate the expected value of the feature being studied. Consequently, we compute the values for each of the 19 MSLR features and 7 similarity scores for our query-document pairs.
@@
Identifying context neurons: After collecting activations and feature labels across a wide range of input sequences (query-document pairs), we start the process of identifying context neurons. Unlike most prior research that employs classification-based probes (MacAvaney et al., 2022; Fan et al., 2021), we utilize ridge regression-based probes to capture the continuous nature of the features under study. We first split the activations and their corresponding labels into training, validate and testing sets (60:20:20).
@@ 4.1.1. MSLR features
For the first set of experiments, we designed sparse probes for each of the selected MSLR features and conducted experiments on each layer of the RankLlama2-7b architecture. All experiments were run with both mean and max activation aggregation over the input sequence. Our findings (Figure 2) are categorized into two groups: (i) Features that exhibit a strong fit in certain layers, (ii) Features that do not correlate to the LLM activations in any layer
Positives: Features that achieve an $R^{2}$ score greater than 0.85 in any particular layer are considered strong positives. This means it is highly likely that the activation has extracted the particular feature within the neuron. We found that seven metrics, namely Min of TF, Min(TF/L), Min of TF*IDF, covered QT Number, covered QT Ratio, Mean(TF/L), and variance of TF*IDF frequently exhibit low Mean Square Errors and high $R^{2}$ scores. This indicates that there exist MLP neurons within the investigated layers that perform feature extractions similar to these MSLR features. Additionally, we observed that the $R^{2}$ scores obtained with mean and max activation aggregation are comparable, except minor exceptions, suggesting either aggregation method is suitable.
Negatives: Certain features failed to achieve an $R^{2}$ score greater than 0.1 in the final layer, often achieving a highly negative $R^{2}$. This indicates that the LLM does not consider these features important in their current form. The 10 features from the MSLR set that fell into this category include: Sum and Max of TF*IDF, Sum(TF/L) and Max(TF/L), Max and variance of TF, stream length and Sum of TF. Thus, while Mean(TF/L) is something the LLM highly seeks, it does not seek for the sum and max of normalized term frequency at all!
@@ 4.1.2. Traditional similarity metrics
Figure 5. Probing RankLlama2-7b and 13b with in-distribution vs out-of-distribution datasets. We witness that most MSLR features when probed for in RankLlama2-7b, show similar performance with both in-distribution and out-of-distribution datasets. This is however not the case with RankLlama2-13b, where certain features like stream length and sum of term frequency show a strong presence in the in-distribution dataset probe, even though they are unlikely features to influence a ranking decision. This suggests overfitting on the MS MARCO dataset in RankLlama2-13b.
@@ 4.1.3. BERT and T5 scores
BERT and T5 are neural models widely used for retrieval and reranking tasks before the advent of LLMs. LLMs are much larger in size compared to BERT models. As a result it is of interest to see if BERT and T5 subnetworks are present within the studied LLM’s activations. We design probes to mine for the cosine distance between the BERT and T5 embeddings of the query and the document. Our observations show that both BERT and T5 obtain moderate $R^{2}$ scores in our probes, reaching $0.7$ and $0.82$ on RankLlama2-7b and 13b respectively. The BERT and T5 scores follow each other across the LLM layers. This indicates that the LLM likely does not encode BERT and T5 subnetworks as-is. We compare our findings to related works on BERT probing in Section 5.

### v3

# Probing Ranking LLMs: A Mechanistic Analysis for Information Retrieval

@@ Abstract.
Transformer networks, particularly those achieving performance comparable to GPT models, are well known for their robust feature extraction abilities. However, the nature of these extracted features and their alignment with human-engineered ones remain largely unexplored. In this work, we investigate the internal mechanisms of state-of-the-art, fine-tuned LLMs for passage reranking. We employ a probing-based analysis to examine neuron activations in ranking LLMs, identifying the presence of known human-engineered and semantic features. Our study spans a broad range of feature categories, including lexical signals, document structure, query-document interactions, and complex semantic representations, to uncover underlying patterns influencing ranking decisions.
Through experiments on four different ranking LLMs, we identify statistical IR features that are prominently encoded in LLM activations, as well as others that are notably missing. Furthermore, we analyze how these models respond to out-of-distribution queries and documents, revealing distinct generalization behaviors. By dissecting the latent representations within LLM activations, we provide an important initial step toward better intepretability and aim to improve both the interpretability and effectiveness of ranking models. Our findings offer crucial insights for developing more transparent and reliable retrieval systems. We release all necessary scripts and code to support further exploration.
Mechanistic Interpretability, Probing, Reranking LLMs, LoRA
journalyear: 2025copyright: ccconference: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR); July 18, 2025; Padua, Italybooktitle: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR) (ICTIR ’25), July 18, 2025, Padua, Italydoi: 10.1145/3731120.3744603isbn: 979-8-4007-1861-8/2025/07ccs: Information systems Learning to rank
@@ 1.1. Contributions and Findings
    We discovered that several human-engineered metrics, such as covered query term number, covered query term ratio, mean of stream length normalized term frequency, and variance of tf*idf, are prominently encoded within LLM activations. In contrast, certain features, including sum of stream length normalized term frequency, max of stream length normalized term frequency, and BM25, show no discernible representation in LLM activations.
@@
    We fine-tuned and probed two additional reranking models, Llama3.1-8b and Pythia-6.9b, resulting in RankLlama3-8b and RankPythia, alongside probing the existing RankLlama2-7b and RankLlama2-13b models. The results revealed broadly consistent outcomes across all models, suggesting that different LLM architectures, when fine-tuned similarly, encode latent features in a comparable way within their activations.
@@
    Finally, we show further evidence on the presence of these encoded features using Shapley values. All scripts, datasets, and fine-tuned ranking LLMs developed and/or used in this study are made publicly available. [^1] https://github.com/TaKneeAa/ProbingRankLlama
@@ 2.1. Mechanistic Interpetability & Sparse Probing
Mechanistic Interpretability is the ambitious goal of gaining an algorithmic-level understanding of any deep neural network’s computations (Räuker et al., 2023). This goal can be achieved in numerous ways, some of them being by studying weights (e.g., weight masking (Csordás et al., 2020; Wortsman et al., 2020), continual learning (De-Arteaga et al., 2019)), by studying individual neurons (e.g., excitement based (Zhou et al., 2014), activation patching (Makelov et al., 2023), gradient-based (Ancona et al., 2019), perturbation and ablation based  (Zhou et al., 2018)), by studying subnetworks (e.g., sparsity based  (Meister et al., 2021), circuit analysis based  (Fiacco et al., 2019)), by studying representations (e.g., tokens  (Li et al., 2021), attention  (Clark et al., 2019), probing  (Belinkov, 2022; Gurnee and Tegmark, 2023; Chen et al., 2023)), by training sparse auto encoders  (Makelov et al., 2024; Kissane et al., 2024; Cunningham et al., 2023), etc. In this study, we focus on studying the activations of individual neurons and also some groups of neurons with the help of probing regressors. To that end, we use concepts from a host of techniques listed above.
@@
Probing aims to determine if a given representation effectively captures a specific type of information, as discussed by Belinkov (Belinkov, 2022). This technique employs transfer learning to test whether embeddings contain information pertinent to a target task. The three essential steps in probing include: (1) obtaining a dataset with examples that exhibit variation in a particular quality of interest, (2) embedding these examples, and (3) training a model on these embeddings to assess if it can learn the quality of interest. This method is versatile as it can utilize any inner representation from any model. However, a limitation of probing methods is that in some cases, a successful probe does not necessarily mean that the probed model actually utilizes that information about the data  (Ravichander et al., 2020). Belinkov (Belinkov, 2022) provides a comprehensive survey on probing methods for large language models, discussing their advantages, disadvantages, and complexities. Gurnee et al.  (Gurnee et al., 2023) further introduce sparse probing, where they mine features of interest over groups of representations up to $k$ in size. We build upon sparse probing in this work.
@@
Table 1. The Ranking effectiveness of various fine-tuned ranking LLMs, in comparision to the performance of a BM25 reranker. We note that while LLMs are effecient for reranking, they are opaque and complex to understand.
@@ 2.2. Inner Interpretability in IR
Most works of interpretability in information retrieval  (Anand et al., 2022) have been model extrinsic (Rahimi et al., 2021; Singh and Anand, 2019; Chowdhury et al., 2023). Among model intrinsic interpretability methods, some works focus on using gradient-based methods to identify important neurons  (Chen et al., 2024; Fernando et al., 2019). Although gradient-based methods give an accurate perspective on the flow of information, they are too fine-grained to give a human-level understanding of the LLM’s inner circuit (Anand et al., 2022). A number of works have attempted to examine the inner workings of neural retrievers to understand if they satisfy IR axioms and/or to spot known features. Fan et al. (Fan et al., 2021) probe fine-tuned BERT models on three different tasks, namely document retrieval, answer retrieval, and passage retrieval, to find out if these different forms of relevance lead to different models. Zhan et al.  (Zhan et al., 2020) study the attention patterns of BERT after fine-tuning on the document ranking task and make note of how BERT dumps duplicate attention weights on high frequency tokens (like periods). Similar to them, Choi et al. (Choi et al., 2022) study attention maps in BERT and report the discovery of the IDF feature within BERT attention patterns. ColBERT investigations (Formal et al., 2021, 2022) study its term-matching mechanism and conclude that it captures a notion of term importance, which is enhanced by fine-tuning. MacAvaney et al.  (MacAvaney et al., 2022) propose ABNIRML, a set of diagnostic probes for neural retrieval models that allows searching for features like writing styles, factuality, sensitivity and word order in models like ColBERT and T5. Parry et al.  (Parry et al., 2024) provide a framework for diagnostic analysis and intervention within Neural Ranking models by generating perturbed inputs followed by activation patching.
These studies, however, have been conducted in the pre-LLM era and hence are much smaller in size. It is unknown if the findings from BERT will carry over to LLMs like Llama, given that BERT is an encoder-only model whereas most modern LLMs are decoder-only models. While most of the above works focus on attention heads and pattern, our work focuses on probing the MLP activation layers, which are now believed to be the primary feature extraction location within the LLM (Gurnee et al., 2023). Parallel to our work, Liu et al. (Liu et al., 2025), through activation patching experiments, identify a progressive, multi-stage process where LLMs extract document and query information in early layers, process relevance according to instructions in middle layers, and use specific attention heads in later layers to generate relevance judgments. Recently Nijasure et al. (Nijasure et al., 2025) conduct a behavioral exploration of LoRA fine-tuned LLMs for Passage Reranking to understand how relevance signals are learned and deployed by them. To the best of our knowledge, our work is one of the first few works towards the study of neuron activations in LLMs for a large scale of statistical and semantic features.
@@ 3. Identifying Context Neurons
Figure 2. Probing for statistical features from the MSLR dataset in RankLlama2-7b model. Here $QT$ stands for Query Term, $TF$ stands for Term Frequency and $\cdot/L$ stands for length normalized. The graph lines indicate the presence of a particular feature along the layers of the LLM. Other features like $Covered\ QT\ Number$, $Covered\ QT\ Ratio$, $Mean(TF/L)$ and $Variance\ TF*IDF$ show increasing prominence from the 1st layer to the last, ultimately playing an important role in ranking decision making. Other MSLR features like $Sum(TF/L)$, $Max(TF/L)$, and $Sum\ TF*IDF$ show negative correlation with RankLlama decision making.
@@
Target Features: The MSLR dataset (Qin and Liu, 2013) provides a collection of features that that can be used for learning to rank experiments. Because we are calculating these features ourselves for some additional documents, many of the MSLR features are not usable in our study – e.g., we do not have anchors, links, URLs, or similar information. Consequently, we aim to search for a subset of these MSLR features within the LLM activations. Specifically, we focus on mining the following 19 features from the MSLR dataset within the LLM (“stream” means the document or passage text): Min TF, Min(TF/L), Min TF*IDF, Covered QT Number, Covered QT Ratio, Mean(TF/L), Variance TF*IDF, BM25, Mean TF*IDF, Variance(TF/L), Mean TF, Variance TF, Sum TF, Max TF, Max(TF/L), Stream length, Sum TF, Max TF*IDF and Sum TF*IDF. We do not claim that this is a definitive list of features universally acknowledged as important for ranking; they are chosen because they are well known.
@@
In addition to MSLR features, we mine for known query-document similarity metrics within the LLM architecture. For this we consider five traditional query-document similarity scores like tf*idf cosine scores, Euclidian score, Manhattan score, KL-divergence score, Jensen-Shannon divergence score as well as two popular semantic relevance metrics like BERT and T5 scores.
@@
Each feature is probed individually: for instance, the probe for stream length is dedicated solely to that feature without searching for any others. Although many of these features are known to be correlated and may share neurons, this study does not include correlation analysis among features.
@@
Probing Datasets: To facilitate probing (Sajjad et al., 2022), we need to curate a dataset of input sequences to study the model. We select query-document pairs from the MS MARCO test set (Nguyen et al., 2016) for this purpose. For each query, we include documents that are highly relevant, highly non-relevant, and of intermediate relevance. Our input sequences for the ranking LLM consist of these query-document pairs. Note that it is important for each probing dataset to be balanced – i.e., contain a uniform number of samples across the range of values of the feature being probed (Gurnee et al., 2023; Belinkov, 2022). For instance, when analyzing the BM25 feature, our dataset included a balanced ratio of query-documents with both low and high BM25 values. Having an unbalanced probing dataset can lead to a biased analysis as the model might have a tendency to overfit the dominant class. For each input sequence, we calculate the expected value of the feature being studied. Consequently, we compute the values for each of the 19 MSLR features and 7 similarity scores for our query-document pairs.
@@
Identifying context neurons: After collecting activations and feature labels for a wide range of input sequences (query-document pairs), we start the process of identifying context neurons. Unlike most prior research that employs classification-based probes (MacAvaney et al., 2022; Fan et al., 2021), we utilize ridge regression-based probes to capture the continuous nature of the features under study. We first split the activations and their corresponding labels into training, validate and testing sets (60:20:20).
@@ 4.1.1. MSLR features
For the first set of experiments, we designed sparse probes for each of the selected MSLR features and conducted experiments on each layer of the RankLlama2-7b architecture. All experiments were run with both mean and max activation aggregation over the input sequence. Our findings (Figure 2) using mean aggregation and are categorized into two groups: (i) Features that exhibit a strong fit in certain layers, (ii) Features that do not correlate to the LLM activations in any layer.
Positives: Features that achieve an $R^{2}$ score greater than 0.85 in any particular layer are considered strong positives. This means it is highly likely that the activation has extracted the particular feature within the neuron. We found that four metrics, namely covered QT Number, covered QT Ratio, Mean(TF/L), and variance of TF*IDF frequently exhibit low Mean Square Errors and high $R^{2}$ scores. This indicates that there exist MLP neurons within the investigated layers that perform feature extractions similar to these MSLR features. Additionally, we observed that the $R^{2}$ scores obtained with mean and max activation aggregation are comparable, except minor exceptions, suggesting either aggregation method is suitable.
Negatives: Certain features failed to achieve an $R^{2}$ score greater than 0.1 in the final layer, often achieving a highly negative $R^{2}$. This indicates that the LLM does not consider these features important in their current form. The 10 features from the MSLR set that fell into this category include: Sum and Max of TF*IDF, Sum(TF/L) and Max(TF/L), Max and variance of TF, stream length and Sum of TF. Thus, while Mean(TF/L) is something the LLM highly seeks, it does not seek for the sum and max of normalized term frequency at all.
@@ 4.1.2. Traditional similarity metrics
Figure 5. Probing the final layers of RankLlama2-7b and 13b with in-distribution vs out-of-distribution datasets. We witness that most MSLR features when probed for in RankLlama2-7b, show similar performance with both in-distribution and out-of-distribution datasets. This is however not the case with RankLlama2-13b, where certain features like stream length and sum of term frequency show a strong presence in the in-distribution dataset probe, even though they are unlikely features to influence a ranking decision. This suggests overfitting on the MS MARCO dataset in RankLlama2-13b.
@@ 4.1.3. BERT and T5 scores
BERT and T5 are neural models widely used for retrieval and reranking tasks before the advent of LLMs. LLMs are much larger in size compared to BERT models. As a result it is of interest to see if BERT and T5 subnetworks are present within the studied LLM’s activations. We design probes to mine for the cosine distance between the BERT and T5 embeddings of the query and the document. Our observations show that both BERT and T5 obtain moderate $R^{2}$ scores in our probes, reaching $0.7$ and $0.82$ on RankLlama2-7b and 13b respectively. The BERT and T5 scores follow each other across the LLM layers. This suggests the LLM likely does not encode BERT and T5 subnetworks as-is. We compare our findings to other work on BERT probing in Section 5.